OpenStudy (anonymous):

User: Convert the equation to the standard form for a hyperbola by completing the square on x and y. y^2-25x^2 + 4y + 50x - 46 = 0 a. (y+4)^2/25-(x-2)2=1 b.(x+2)^2/25-(y-1)2=1 c. (x-1)2-(x=2)^2/25=1 d.(x=2)^2/25-(x-1)2=1

OpenStudy (anonymous):

\[y^2+4y-25(x^2-2)=46\] \[(y+2)^2-25(x-1)^2=46+4-25=25\] \[\frac{(y+2)^2}{25}-(x-1)^2=1\]